What is the magnitude of gravitational force due to the interaction of 1000g and 2000kg mass particle at 2 cm distance between them?


1
Expert's answer
2021-11-15T10:32:55-0500
"F=\dfrac{Gm_1m_2}{d^2},""F=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times1\ kg\times2000\ kg}{(0.02\ m)^2}=3.33\times10^{-4}\ N."